OpenStudy (anonymous):

Point Slope Form of (2,-5) and (4,3)

OpenStudy (triciaal):

slope, m = change in y divided by corresponding change in x so y2-y1/x2 - x1 for 2 given points (x1,y1) and (x2,Y2) on a line for any point on the line (x,y) the slope will be the same y - y1 = m(x-x1) same as y-y2= m(x - x2) for this question the 2 given points (x1,y1) and (x2,y2) are (2,-5) and (4,3) so plug the numbers to find m then use this value of m with one of the points to write the general equation to include point (x,y)